At 60oC the value of Kw is 90.5 x 10-4. Considering this, what is the calculated value for the pH of solution made by dissolving 1.00g of sodium hydroxide in enough water to make 500ml of solution at this temperature?
1
Expert's answer
2020-07-16T07:00:05-0400

There is an error in the formulation of the problem. Kw cannot be equal to 10-4, it must be equal to 10-14.


nNaOH = 1(g)/40(g/mol) = 0.025 mol

CNaOH = n/V = 0.25 mol / 0.5 l = 0.05 mol/l

C[H+] = Kw / CNaOH = 90.5•10-14/0.05 = 1.81•10-11 mol/l

pH = -lgC[H+] = 10,74


Answer is pH=10,74